# Explicit Motion Handling and Interactive Prompting for Video Camouflaged Object Detection

### v1

@@ Abstract
Camouflage poses challenges in distinguishing a static target, whereas any movement of the target can break this disguise. Existing video camouflaged object detection (VCOD) approaches take noisy motion estimation as input or model motion implicitly, restricting detection performance in complex dynamic scenes. In this paper, we propose a novel Explicit Motion handling and Interactive Prompting framework for VCOD, dubbed EMIP, which handles motion cues explicitly using a frozen pre-trained optical flow fundamental model. EMIP is characterized by a two-stream architecture for simultaneously conducting camouflaged segmentation and optical flow estimation. Interactions across the dual streams are realized in an interactive prompting way that is inspired by emerging visual prompt learning. Two learnable modules, i.e. the camouflaged feeder and motion collector, are designed to incorporate segmentation-to-motion and motion-to-segmentation prompts, respectively, and enhance outputs of the both streams. The prompt fed to the motion stream is learned by supervising optical flow in a self-supervised manner. Furthermore, we show that long-term historical information can also be incorporated as a prompt into EMIP and achieve more robust results with temporal consistency. Experimental results demonstrate that our EMIP achieves new state-of-the-art records on popular VCOD benchmarks. The code will be publicly available.
1 Introduction
@@
Camouflaged object detection (COD) aims at detecting and segmenting those hidden objects that exhibit high intrinsic similarity to their backgrounds. The inherent complexity of distinguishing camouflaged objects from their surroundings poses unique challenges when compared to general object detection [?] and salient object detection [?]. Recently, it has attracted interest of many researchers and facilities broad applications, e.g., medical image segmentation [??] and industrial inspection [??]. While significant progress has been made in image-based [???] and video-based [???] COD tasks, there still remains substantial room for development due to intrinsic difficulty of the tasks.
@@
Figure 1: Different strategies of motion handling in VCOD: (a) Directly feed optical flow maps [??]; (b) Learn implicit motion cues and subsequently utilize them for mask decoding [?]; (c) The proposed interactive prompting paradigm handles motion cues explicitly using a pre-trained optical flow model, and simultaneously conduct optical flow estimation and camouflaged object segmentation. The fire/snowflake symbols denote that most of the model parameters are learnable/frozen in the proposed scheme.
@@
Image-based COD approaches identify camouflaged targets using a static image. Typically, these targets exhibit strong visual resemblances to their backgrounds in terms of texture, color, or edges, posing a challenge for detection using appearance or geometric cues alone. Consequently, motion cues have been investigated in previous works [??] for video camouflaged object detection (VCOD) task. Fig. 1 summarizes the strategies adopted in handling motion cues by these previous works. As presented in Fig. 1 (a), off-the-shelf motion estimators (e.g., [??]) are directly employed for generating offline optical flow, serving as motion cues for identifying camouflaged objects. Notably, offline motion estimation, particularly in camouflaged scenarios, poses significant challenges for common optical flow estimators, often leading to noisy and inaccurate output optical flow. Such erroneous input could misguide detectors and thereby hinder overall performance. To address this issue, SLT-Net [?] proposes a distinct approach by implicit motion modeling in an online fashion (Fig. 1 (b)). While the overall motion modeling part is learnable, learning reliable motion of camouflaged objects from limited VCOD data may be intractable compared to using extensive training data for training common optical flow estimators [??]. Besides, due to the implicit nature of SLT-Net, no explicit regularization or evaluation can be adhered to the motion part, further making the learned motion less effective and reliable. Also, we believe that beyond motion, appearance cues are an important factor conducive to detection as well, since objects could be motion-free or still.
@@
In this paper, we propose a novel Explicit Motion handling and Interactive Prompting framework for VCOD (Fig. 1 (c)), dubbed EMIP, which handles motion cues explicitly by freezing the upstream optical flow model mostly. EMIP is characterized by a two-stream architecture for simultaneously addressing camouflaged segmentation and optical flow estimation. Inspired by emerging advances of the visual prompt learning [???], we design an interactive prompting scheme to achieve interactions across the dual streams, as shown in Fig. 1 (c). Two modules, namely the camouflage feeder and motion collector, are designed to incorporate segmentation-to-motion and motion-to-segmentation prompts, respectively, and enhance outputs of the dual streams. Due to the absence of authentic optical flow for current VCOD datasets, we learn the prompt fed to the motion stream by supervising optical flow in a self-supervised manner. We also propose a long-term variant of EMIP by formulating historical features into the prompt to mitigate short-term prediction errors and further improve the accuracy. Benefiting from the above elaborate designs, the proposed EMIP  effectively leverages noise-robust motion to detect and segment those video camouflaged objects. The main contributions are summarized below:
@@
  • We also propose a long-term variant of EMIP by formulating historical features into the prompt to mitigate short-term prediction errors.
@@
  • EMIP together with its long-term variant, achieve new state-of-the-art records and outperform previous models by notable margins ($\sim$17.0%/5.5average improvement on F-measure/S-measure over the previous best model SLT-Net).
Figure 2: Overall architecture of the proposed EMIP, which consists of two separate streams: explicit motion modeling stream (upper) and object segmentation stream (lower). We use GMFlow  [?] as the fundamental model to handle motion cues. With the camouflage feeder and motion collector, segmentation and motion prompts are injected into each task-specific stream to compensate essential information. The fire/snowflake symbols indicate that the model parameters in this part or block are designated as learnable/frozen.
2 Related Work
@@
Image-based COD. Methods in this type aim to discern camouflaged objects from a single RGB image. Early COD methods [??] relied on hand-crafted features to find targets hidden in the background. However, with the advent of deep learning, COD methods have undergone substantial advancement in recent years. Inspired by natural predatory behavior, SINet-V2 [?] and PFNet  [?] employed a coarse-to-fine strategy. They first generated a preliminary location map for camouflaged objects and then refined it for segmentation. To enhance performance, several studies integrated auxiliary task into a joint learning framework. MGL [?] combined classification or boundary detection task with COD. ZoomNet [?] employed a zooming in and out strategy to the original inputs and processed appearance features at three different scales. Jia et al. [?] proposed the SegMaR framework, an iterative refinement approach designed to locate, magnify, and detect camouflaged objects. HitNet [?] elevated low-resolution representations by leveraging high-resolution features in an iterative feedback loop, effectively mitigating challenges such as edge blurring and detail degradation.
@@
Video-based COD. For the VCOD task, motion cues are crucial to camouflaged object detection. Bidau et al. [?] introduced a method by approximating various motion models derived from dense optical flow. Lamdouar et al.[?] introduced a video registration and segmentation network to detect camouflaged objects, employing optical flow and a difference image as inputs. However, the utilization of inaccurate optical flow may result in accumulative errors in mask prediction. To address this challenge, Cheng et al. [?] proposed a two-stage model that implicitly models and leverages motion information. Different from the above approaches, we propose to handle motion cues in an explicit way as aforementioned, and a two-stream architecture is designed to conduct both optical flow estimation and segmentation.
@@
Visual Prompt Learning. Recently, prompt learning has emerged as a new paradigm that has significantly enhanced the performance of natural language processing (NLP) tasks [?]. Besides, the prompting paradigm has been adopted in many computer vision tasks [??]. Work [?] modified transformer layers by introducing memory tokens, constituting a set of learnable embedding vectors. VPT [?] employed a similar strategy by applying learnable embedding vectors to transformer encoders, achieving noteworthy performance across various downstream recognition tasks. Based on this idea, ViPT [?] integrated modality-complementary visual prompts for task-oriented multi-modal tracking. The popular Segment Anything Model (SAM) [?] integrates various visual prompts like points, boxes, or masks to achieve tailored object segmentation and exhibits decent zero-shot generalization. Previous researches have mainly focused on specific tasks like classification, tracking, or segmentation. In this paper, inspired by [?], we treat intermediate features from one stream as a prompt for injecting complementary information to the other, and propose an interactive prompting scheme for the VCOD task.
3 Methodology
@@
In this work, we propose an end-to-end trainable network EMIP to jointly optimize camouflaged object detection and optical flow estimation. We input an adjacent frame pair $(I_{t},I_{t-1})$ of a video to EMIP, and output a binary segmentation mask of the reference frame $I_{t}$, together with optical flow estimation $\boldsymbol{V}$. The overall architecture of our EMIP is illustrated in Fig. 2. Note that, for object segmentation stream, a set of features $\{f_{n}^{i}\in\mathbb{R}^{H/{2^{i+1}}\times W/{2^{i+1}}\times C_{i}},n\in\{t,t-1\},i=1,...,4\}$ with different scales are extracted from a vision transformer backbone (i.e., PVTv2-B5 [?]), the same as SLT-Net [?], for fair comparison. $W$, $H$, and $C$ represent the width, height, and channel number, respectively. Similar to [?], we adopt the top-three features ($f_{t}^{2},f_{t}^{3},f_{t}^{4}$) for appearance modeling, and discard the first-layer feature $f_{t}^{1}$.
3.1 Fundamental Model for Motion Modeling
@@
To achieve more effective integration of motion information for the VCOD task, we select GMFlow [?] as our fundamental model, which is trained on $\sim$50k frame pairs for optical flow estimation. As shown in Fig. 2, GMFlow can be delineated into two integral components, i.e., a CNN encoder and a transformer decoder. It initially employs the CNN encoder to capture low-level features such as edges, colors, and textures. Subsequently, the transformer decoder, composed of a sequence of self- and cross-attention layers, predicts optical flow map $\boldsymbol{V}\in\mathbb{R}^{H\times W\times 2}$ and matching distribution $\boldsymbol{M}\in\mathbb{R}^{H\times W\times H\times W}$ simultaneously. $\boldsymbol{V}$ is derived from $\boldsymbol{M}$ through operations such as pixel-grid sampling. Here, we flatten the first two dimensions of $\boldsymbol{M}$ to obtain $\boldsymbol{G}\in\mathbb{R}^{H\times W\times HW}$, and employ $\boldsymbol{G}$ as the motion prompt for interacting with segmentation features, leveraging its detailed pixel-to-pixel matching information. Additionally, the flow map $\boldsymbol{V}$ is utilized to reconstruct frame $I_{t}$ from $I_{t-1}$, leading to the computation of a self-supervised loss. For more details about our fundamental motion estimation model, we refer readers to GMFlow [?].
3.2 Segmentation-to-Motion Prompt

### v2

@@ Abstract
Camouflage poses notable challenges in distinguishing a static target, as it usually blends seamlessly with the background. However, any movement by the target can disrupt this disguise, making it detectable. Existing video camouflaged object detection (VCOD) approaches take noisy motion estimation as input or model motion implicitly, restricting detection performance in complex dynamic scenes. In this paper, we propose a novel Explicit Motion handling and Interactive Prompting framework for VCOD, dubbed EMIP, which handles motion cues explicitly using a frozen pre-trained optical flow fundamental model. EMIP is characterized by a two-stream architecture for simultaneously conducting camouflaged segmentation and optical flow estimation. Interactions across the dual streams are realized in an interactive prompting way that is inspired by emerging visual prompt learning. Two learnable modules, i.e. the camouflaged feeder and motion collector, are designed to incorporate segmentation-to-motion and motion-to-segmentation prompts, respectively, and enhance outputs of the both streams. The prompt fed to the motion stream is learned by supervising optical flow in a self-supervised manner. Furthermore, we show that long-term historical information can also be incorporated as a prompt into EMIP and achieve more robust results with temporal consistency. By leveraging promoting techniques based on EMIP, the proposed long-term model EMIP incurs lower training cost with only 8.5M trainable parameters (less than 8% of the total model parameters). Experimental results demonstrate that both EMIP and EMIP set new state-of-the-art records on popular VCOD benchmarks. Additionally, comparative evaluations against other video segmentation models on a wider range of video segmentation tasks demonstrate the robustness and superior generalization capabilities of EMIP. Our code is made publicly available at https://github.com/zhangxin06/EMIP.
Index Terms:
Video camouflaged object detection, explicit motion modeling, interactive prompting, semantic segmentation, deep learning.
I Introduction
@@
Camouflaged object detection (COD) aims at detecting and segmenting those hidden objects that exhibit high intrinsic similarity to their backgrounds. The inherent complexity of distinguishing camouflaged objects from their surroundings poses unique challenges when compared to general object detection [1] and salient object detection (SOD) [2]. Recently, it has attracted interest of many researchers and facilities broad applications, e.g., medical image segmentation [3, 4] and industrial inspection [5, 6]. While significant progress has been made in image-based [7, 8, 9] and video-based [10, 11, 12] COD tasks, there still remains substantial room for development due to intrinsic difficulty of the tasks.
@@
Figure 1: Different strategies of motion handling in VCOD: (a) Directly feed optical flow maps [11, 13]; (b) Learn implicit motion cues and subsequently utilize them for mask decoding [10]; (c) The proposed interactive prompting paradigm handles motion cues explicitly using a pre-trained optical flow model, and simultaneously conduct optical flow estimation and camouflaged object segmentation. The fire/snowflake symbols denote that most of the model parameters are learnable/frozen in the proposed scheme.
@@
Image-based COD approaches identify camouflaged targets using a static image. Typically, these targets exhibit strong visual resemblances to their backgrounds in terms of texture, color, or edges, posing a challenge for detection using appearance or geometric cues alone. Consequently, motion cues have been investigated in previous works [11, 13] for video camouflaged object detection (VCOD) task. Fig. 1 summarizes the strategies adopted in handling motion cues by these previous works. As presented in Fig. 1 (a), off-the-shelf motion estimators (e.g., [14, 15]) are directly employed for generating offline optical flow, serving as motion cues for identifying camouflaged objects. Notably, offline motion estimation, particularly in camouflaged scenarios, poses significant challenges for common optical flow estimators, often leading to noisy and inaccurate output optical flow. Such erroneous input could misguide detectors and thereby hinder overall performance. To address this issue, SLT-Net [10] proposes a distinct approach by implicit motion modeling in an online fashion (Fig. 1 (b)). While the overall motion modeling part is learnable, learning reliable motion of camouflaged objects from limited VCOD data may be intractable compared to using extensive training data for training common optical flow estimators [14, 16]. Besides, due to the implicit nature of SLT-Net, no explicit regularization or evaluation can be adhered to the motion part, further making the learned motion less effective and reliable. Also, we believe that beyond motion, appearance cues are an important factor conducive to detection as well, since objects could be motion-free or still.
@@
In this paper, we propose a novel Explicit Motion handling and Interactive Prompting framework for VCOD (Fig. 1 (c)), dubbed EMIP, which handles motion cues explicitly by freezing the upstream optical flow model mostly. EMIP is characterized by a two-stream architecture for simultaneously addressing camouflaged segmentation and optical flow estimation. Inspired by emerging advances of the visual prompt learning [17, 18, 19], we design an interactive prompting scheme to achieve interactions across the dual streams, as shown in Fig. 1 (c). Two modules, namely the camouflage feeder and motion collector, are designed to incorporate segmentation-to-motion and motion-to-segmentation prompts, respectively, and enhance outputs of the dual streams. Due to the absence of authentic optical flow for current VCOD datasets, we learn the prompt fed to the motion stream by supervising optical flow in a self-supervised manner. We also propose a long-term variant of EMIP by formulating historical features into the prompt to mitigate short-term prediction errors and further improve the accuracy. Benefiting from the above elaborate designs, the proposed EMIP  effectively leverages noise-robust motion to detect and segment those video camouflaged objects. The main contributions are summarized below:
@@
  • We also propose a long-term variant of EMIP, dubbed EMIP, by formulating historical features into the prompt to mitigate short-term prediction errors.
@@
  • EMIP together with its long-term variant EMIP, achieve new state-of-the-art records and outperform previous models by notable margins ($\sim$17.0%/5.5average improvement from EMIP on F-measure/S-measure over the previous cutting-edge model SLT-Net).
The remainder of the paper is organized as follows: Section II discusses related work on image and video-based COD, closely related video object segmentation, salient object detection, and visual prompt learning. Section III describes the proposed models in detail. Section IV includes experimental results, comparisons, and analyses. Finally, conclusions are drawn in Section V.
II Related Work
II-A Image-based COD
@@
Methods in this type aim to discern camouflaged objects from a single RGB image. Early COD methods [20, 21] relied on hand-crafted features to find targets hidden in the background. Li et al. [22] proposed a texture guided weighted voting strategy to detect camouflaged objects. Then they further introduced a fusion framework [23] to address camouflaged problems in the wavelet domain. And more, Garcia [24] review the background subtraction methods and outlook future direction. However, with the advent of deep learning, COD methods have undergone substantial advancement in recent years. Inspired by natural predatory behavior, SINet-V2 [7] and PFNet  [25] employed a coarse-to-fine strategy. They first generated a preliminary location map for camouflaged objects and then refined it for segmentation. To enhance performance, several studies integrated auxiliary task into a joint learning framework. MGL [26] combined classification or boundary detection task with COD. Liu et al. [27] investigated the part-object relationship to discover camouflaged objects. ZoomNet [28] employed a zooming in and out strategy to the original inputs and processed appearance features at three different scales. Then Pang et al. extended ZoomNet to ZoomNeXt [29], which simultaneously addresses image and video camouflaged object detection. Jia et al. [30] proposed the SegMaR framework, an iterative refinement approach designed to locate, magnify, and detect camouflaged objects. Ji et al. [9] designed a two-branch framework to encode the context and texture of camouflage objects under gradient supervision. Huang et al. [31] designed progressively neighboring token enhanced decoder to exploit imperceptible cues for detecting camouflaged objects. Zhang et al. [32] introduced predictive uncertainty estimation framework to address model and data uncertainty simultaneously in camouflaged scenes . HitNet [33] elevated low-resolution representations by leveraging high-resolution features in an iterative feedback loop, effectively mitigating challenges such as edge blurring and detail degradation. Methods [34, 35] designed frameworks to mine the subtle cues of camouflaged objects in the frequency domain. Yao et al. [36] proposed a hierarchical graph interaction network to refine ambiguous regions. Hao et al. [37] proposed a simple yet effective general architecture for both COD and SOD tasks.
II-B Video-based COD
@@
For the VCOD task, motion cues are crucial to camouflaged object detection. Bidau et al. [38] introduced a method by approximating various motion models derived from dense optical flow. Zhang et al. [39] proposed a camouflage modeling strategy and fused it with discriminative modeling in a Bayesian framework for moving object detection. Lamdouar et al.[11] introduced a video registration and segmentation network to detect camouflaged objects, employing optical flow and a difference image as inputs. However, the utilization of inaccurate optical flow may result in accumulative errors in mask prediction. To address this challenge, Cheng et al. [10] proposed a two-stage model that implicitly models and leverages motion information. Subsequently, to eliminate inaccuracies stemming from implicit motion modeling in SLT-Net [10], Hui et al. [40] introduced a motion-induced consistency preserving approach between frames with a feature pyramid framework. Lu et al. [41] introduced a weakly-supervised framework for VCOD. Hui et al. [42] leveraged temporal and spatial relationships between frames to generate prompts for SAM. Different from the above approaches, we propose to handle motion cues in an explicit way as aforementioned, and a two-stream architecture is designed to conduct both optical flow estimation and segmentation.
II-C Video Object Segmentation (VOS) & Video Salient Object Detection (VSOD)
VCOD can be seen as a specialized task of video object segmentation (VOS) that segments objects across consecutive video frames. Mei et al. [43] proposed a transformer-based framework to leverage temporal and spatial relationships across frames. To increase the detection speed, Park et al. [44] introduced a novel network capable of dynamically selecting mask generation methods, either by reusing features from prior frames or processing the entire network. To incorporate these advantages, memory-based networks [45, 46] have also been explored to better utilize historical information. These VOS networks use the current frame to query a memory bank storing historical features and corresponding object masks from past frames.
On the other hand, the objective of video salient object detection (VSOD) stands in direct contrast to VCOD. VSOD focuses on locating and segmenting the most prominent objects from sequences. Chen et al. [47] utilized spatial and temporal cues along with local constraints to achieve global saliency optimization. Li et al. [48] proposed a motion-based bilateral network for background estimation, and the background estimation results are then merged with instance embeddings into a graph, where edges connect pixels across different frames for multi-frame reasoning. Song et al. [49] proposed a novel recurrent network to extract multi-scale spatial features, which are then concatenated and fed into an extended deep bidirectional ConvLSTM to learn spatiotemporal information. Cong et al. [50] designed a two-stage method: first, obtaining the spatial saliency of each frame through sparsity-based reconstruction, and then capturing the sequential correspondence in the temporal space via progressive sparsity-based propagation. Xu et al. [51] proposed a novel method for modeling motion energy based on four aspects: gradient flow field, motion direction, motion magnitude, and the spatial gradients of video frames. Yan et al. [52]  et al. leveraged optical flow estimation to generate pseudo-labels for some unannotated frames in the dataset, and further enhances the spatio-temporal correlation between video frames using Non-local [53] and ConvGRU [54]. Chen et al. [55] integrated a lightweight temporal model into the spatial branch, coarsely locating spatial saliency regions associated with highly confident salient motion. Simultaneously, the spatial branch itself can iteratively optimize the temporal model in a multi-scale manner. Ji et al. [56] proposed a full-duplex strategy to obtain more stable consistent features. Zhao et al. [57] introduced a space-time memory-based network and leveraged high-level features to refine low-level details.
Compared to existing VOS and VSOD methods, our model excels in discerning subtle differences between camouflaged objects and their highly similar surroundings by simultaneously integrating camouflage properties with explicit motion modeling information.
II-D Visual Prompt Learning
@@
Recently, prompt learning has emerged as a new paradigm that has significantly enhanced the performance of natural language processing (NLP) tasks [58]. Besides, the prompting paradigm has been adopted in many computer vision tasks [17, 18]. Work [17] modified transformer layers by introducing memory tokens, constituting a set of learnable embedding vectors. VPT [18] employed a similar strategy by applying learnable embedding vectors to transformer encoders, achieving noteworthy performance across various downstream recognition tasks. Based on this idea, ViPT [19] integrated modality-complementary visual prompts for task-oriented multi-modal tracking. The popular Segment Anything Model (SAM) [59] integrated various visual prompts like points, boxes, or masks to achieve tailored object segmentation and exhibited decent zero-shot generalization. Previous researches have mainly focused on specific tasks like classification, tracking, or segmentation. In this paper, inspired by [19], we treat intermediate features from one stream as a prompt for injecting complementary information to the other, and propose an interactive prompting scheme for the VCOD task.
III Methodology
@@
In this work, we propose an end-to-end trainable network EMIP to jointly optimize camouflaged object detection and optical flow estimation. We input an adjacent frame pair $(I_{t},I_{t-1})$ of a video to EMIP, and output a binary segmentation mask of the reference frame $I_{t}$, together with optical flow estimation $\boldsymbol{V}$. The overall architecture of our EMIP is illustrated in Fig. 2. Note that, for object segmentation stream, a set of features $\{f_{n}^{i}\in\mathbb{R}^{H/{2^{i+1}}\times W/{2^{i+1}}\times C_{i}},n\in\{t,t-1\},i=1,...,4\}$ with different scales are extracted from a vision transformer backbone (i.e., PVTv2-B5 [60]), the same as SLT-Net [10], for fair comparison. $W$, $H$, and $C$ represent the width, height, and channel number, respectively. Similar to [61], we adopt the top-three features ($f_{t}^{2},f_{t}^{3},f_{t}^{4}$) for appearance modeling, and discard the first-layer feature $f_{t}^{1}$.
Figure 2: Overall architecture of the proposed EMIP, which consists of two separate streams: explicit motion modeling stream (upper) and object segmentation stream (lower). We use GMFlow [16] as the fundamental model to handle motion cues. With the camouflage feeder and motion collector, segmentation and motion prompts are injected into each task-specific stream to compensate essential information. The fire/snowflake symbols indicate that the model parameters in this part or block are designated as learnable/frozen.
III-A Fundamental Model for Motion Modeling
@@
To achieve more effective integration of motion information for the VCOD task, we select GMFlow [16] as our fundamental model, which is trained on $\sim$50k frame pairs for optical flow estimation. As shown in Fig. 2, GMFlow can be delineated into two integral components, i.e., a CNN encoder and a transformer decoder. It initially employs the CNN encoder to capture low-level features such as edges, colors, and textures. Subsequently, the transformer decoder, composed of a sequence of self- and cross-attention layers, predicts optical flow map $\boldsymbol{V}\in\mathbb{R}^{H/8\times W/8\times 2}$ and matching distribution $\boldsymbol{M}\in\mathbb{R}^{H/8\times W/8\times H/8\times W/8}$ simultaneously. $\boldsymbol{V}$ is derived from $\boldsymbol{M}$ through operations such as pixel-grid sampling. Here, we flatten the first two dimensions of $\boldsymbol{M}$ to obtain $\boldsymbol{G}\in\mathbb{R}^{H/8\times W/8\times HW/64}$, and employ $\boldsymbol{G}$ as the motion prompt for interacting with segmentation features, leveraging its detailed pixel-to-pixel matching information. Additionally, the optical flow map $\boldsymbol{V}$ is utilized to reconstruct frame $I_{t}$ from $I_{t-1}$, leading to the computation of a self-supervised loss. For more details about our fundamental motion estimation model, we refer readers to GMFlow [16].
III-B Segmentation-to-Motion Prompt